# Sliced Wasserstein Discrepancy in Disentangling Representation and Adaptation Networks for Unsupervised Domain Adaptation

### v1

@@ Abstract
This paper introduces DRANet-SWD, an extension of existing work that disentangles content and style representations of images for unsupervised domain adaptation (UDA). The approach builds upon DRANet by incorporating the sliced Wasserstein discrepancy (SWD) as a style loss instead of the traditional Gram matrix loss. The potential advantages of SWD over the Gram matrix loss for capturing style variations in domain adaptation are investigated. Experiments using digit classification datasets and driving scenario segmentation validate the method, demonstrating that DRANet-SWD enhances performance. Results indicate that SWD provides a more robust statistical comparison of feature distributions, leading to better style adaptation. These findings highlight the effectiveness of SWD in refining feature alignment and improving domain adaptation tasks across these benchmarks. Our code can be found here.
@@ III-A Disentangling Representation and Adaptation Networks
The DRANet-SWD architecture is built off of the previous work from DRANet[4]. DRANet consists of multiple networks which work to take two or more image datasets and gain an understanding of the image’s content and style. DRANet uses these content and style representations in order to adapt the content of one image into the domain of another using the other domain’s style representation. The architecture of DRANet is shown in Figure 1. The networks that make up DRANet are an encoder $E$, separator $S$, generator $G$, two discriminators $D_{X}$, $D_{Y}$, and a perceptual network $P$.
@@ V-A Digit Classification
The digit classification results are shown in Table II and the associated images are shown in Figure 2. DRANet-SWD shows an improvement in most adaptations with exceptions being MNIST to MNIST-M and SVHN to MNIST. DRANet-SWD shows a great improvement for adapting MNIST to USPS and vice versa. These datasets are both grayscale with differing resolutions, USPS being 16x16 and MNIST being 28x28, as well as differing stroke styles and thicknesses. These datasets see a valuable improvement in classification accuracy likely due to SWD more effectively capturing the textural style between images with differing resolutions. SWD functioning with probability distributions may generalize better in this case where standard feature alignment has more difficulty bridging this gap. DRANet-SWD nearly matches the performance of DRANet for adapting MNIST to MNIST-M and vice versa. It does seem to misinterpret a style rarely for MNIST to MNIST-M but most styles seem to have improved fidelity compared to DRANet. The content of the images produced by DRANet-SWD seem much clearer especially for MNIST-M to MNIST with several examples showing DRANet-SWD succeeds where DRANet fails. The domain transfer tasks involving SVHN show the limitations of the content and style representation approach. Other latent space approaches for this specific dataset such as triplet loss networks [30] have superior performance. In the SVHN dataset, multiple numbers are present confusing content and style representations which severely impacts performance. Interestingly when examining Figure 2(c) DRANet-SWD has partially and poorly adapted some digits, while DRANet was unable. In order to achieve the results shown in Table II for this MNIST to SVHN and vice versa, a more expressive classifier was required than was used for the other digit classification tasks.
@@ V-B Driving Scene Segmentation
The results for the GTA5 to Cityscapes segmentation are shown in Table I and the segmented images shown in Figure 3. The results compare DRANet and DRANet-SWD with and without the further training for the segmentation network. Overall DRANet-SWD outperforms DRANet achieving slightly higher mean Intersection-over-Union (mIoU) and demonstrates higher accuracy in many categories. Interestingly DRANet achieves better accuracy in several categories such as road, sidewalk, car and truck without pretraining. However, DRANet-SWD shows an improvement when classifying smaller or less frequently occurring objects, suggesting that the SWD mechanism improves robustness to rare categories. Its important to note that these results under perform in comparison to the results reported by Lee et al.[4] suggesting differing training setup. An avenue for further experimentation include additional hyperparameter tuning including batch size, SWD projection number etc to ensure similar results as previously reported.
These results were obtained using a fixed random seed (5688), as provided in the original DRANet GitHub repository.
VI Conclusion
In this paper, we introduced DRANet-SWD, a disentangled representation learning framework for unsupervised domain adaptation that replaces the traditional Gram matrix style loss with sliced Wasserstein discrepancy (SWD). Overall, our implementation of DRANet, from the official GitHub source, did not achieve the originally reported results. However, our proposed method, DRANet-SWD, consistently outperforms this implementation, demonstrating its effectiveness. By leveraging SWD, our model better captures the underlying statistical properties of style representations, leading to domain adaptation performance that is better or equivalent to the existing method. Our experiments demonstrate that DRANet-SWD enhances visual fidelity and adaptation effectiveness across both classification and segmentation tasks, highlighting the advantages of SWD over conventional style transfer methods. We also discuss and showcase the datasets where the model struggles. These results reinforce the importance of distribution-based metrics in representation learning and provide a strong foundation for further advancements in domain adaptation. Additional avenues for improvement include analyzing existing or developing new representation strategies for adapting difficult problems such as adapting MNIST and SVHN unsupervised.

### v2

@@ Abstract
This paper introduces DRANet-SWD as a novel complete pipeline for disentangling content and style representations of images for unsupervised domain adaptation (UDA). The approach builds upon DRANet by incorporating the sliced Wasserstein discrepancy (SWD) as a style loss instead of the traditional Gram matrix loss. The potential advantages of SWD over the Gram matrix loss for capturing style variations in domain adaptation are investigated. Experiments using digit classification datasets and driving scenario segmentation validate the method, demonstrating that DRANet-SWD enhances performance. Results indicate that SWD provides a more robust statistical comparison of feature distributions, leading to better style adaptation. These findings highlight the effectiveness of SWD in refining feature alignment and improving domain adaptation tasks across these benchmarks. Our code can be found here.
@@ III-A Disentangling Representation and Adaptation Networks
The DRANet-SWD architecture is built off of the previous work from DRANet[4]. DRANet consists of multiple networks which work to take two or more image datasets and gain an understanding of the image’s content and style. DRANet uses these content and style representations in order to adapt the content of one image into the domain of another using the other domain’s style representation. The architecture of DRANet is shown in Fig 1. The networks that make up DRANet are an encoder $E$, separator $S$, generator $G$, two discriminators $D_{X}$, $D_{Y}$, and a perceptual network $P$.
@@ V-A Digit Classification
The digit classification results are shown in Table II and the associated images are shown in Fig 2. DRANet-SWD shows an improvement in most adaptations with exceptions being MNIST to MNIST-M and SVHN to MNIST. DRANet-SWD shows a great improvement for adapting MNIST to USPS and vice versa. These datasets are both grayscale with differing resolutions, USPS being 16x16 and MNIST being 28x28, as well as differing stroke styles and thicknesses. These datasets see a valuable improvement in classification accuracy likely due to SWD more effectively capturing the textural style between images with differing resolutions. SWD functioning with probability distributions may generalize better in this case where standard feature alignment has more difficulty bridging this gap. DRANet-SWD nearly matches the performance of DRANet for adapting MNIST to MNIST-M and vice versa. It does seem to misinterpret a style rarely for MNIST to MNIST-M but most styles seem to have improved fidelity compared to DRANet. The content of the images produced by DRANet-SWD seem much clearer especially for MNIST-M to MNIST with several examples showing DRANet-SWD succeeds where DRANet fails. The domain transfer tasks involving SVHN show the limitations of the content and style representation approach. Other latent space approaches for this specific dataset such as triplet loss networks [30] have superior performance. In the SVHN dataset, multiple numbers are present confusing content and style representations which severely impacts performance. Interestingly when examining Fig 2(c) DRANet-SWD has partially and poorly adapted some digits, while DRANet was unable. In order to achieve the results shown in Table II for this MNIST to SVHN and vice versa, a more expressive classifier was required than was used for the other digit classification tasks.
@@ V-B Driving Scene Segmentation
The results for the GTA5 to Cityscapes segmentation are shown in Table I and the segmented images shown in Fig 3. The results compare DRANet and DRANet-SWD with and without the further training for the segmentation network. Overall DRANet-SWD marginally outperforms DRANet achieving slightly higher mean Intersection-over-Union (mIoU) and demonstrates higher accuracy in several categories. Interestingly, DRANet achieves better accuracy in several categories such as road, sidewalk, car and truck without pretraining. However, DRANet-SWD shows an improvement when classifying smaller or less frequently occurring objects, suggesting that the SWD mechanism improves robustness to rare categories. Its important to note that these results under perform in comparison to the results reported by Lee et al.[4] suggesting differing training setups. An avenue for further experimentation include additional hyperparameter tuning including batch size, SWD projection number etc to ensure similar results as previously reported.
These results were obtained using a fixed random seed (5688), as provided in the DRANet GitHub repository.
VI Conclusions
In this paper, we introduced DRANet-SWD, a disentangled representation learning framework for unsupervised domain adaptation that replaces the traditional Gram matrix style loss with sliced Wasserstein discrepancy (SWD). While our independent implementation of DRANet did not achieve the originally reported results, the incorporation of SWD into this method in the proposed pipeline consistently improved results. Specifically, our model better captures the underlying statistical properties of style representations by leveraging SWD, leading to domain adaptation performance that is better or equivalent to the existing method. Our experiments demonstrate that DRANet-SWD enhances visual fidelity and adaptation effectiveness across both classification and segmentation tasks, highlighting the advantages of SWD over traditional style transfer methods. We also discuss and showcase the datasets where the model struggles. These results reinforce the importance of distribution-based metrics in representation learning and provide a strong foundation for further advancements in domain adaptation. Additional avenues for improvement include analyzing existing or developing new representation strategies for adapting difficult problems such as adapting MNIST and SVHN unsupervised.